OpenStudy (anonymous):

I am confused on this problem the number of years a person born in the united states is expected to live can approximated by the equation y= 0.189x +70.8 where x is the number of years since 1970 solve the equation for x use this new equation to determine in which year the approximate life expectancy will be 80.2 years.

OpenStudy (phi):

solve the equation for x y= 0.189x +70.8 what do you get if you add -70.8 to both sides ?

OpenStudy (anonymous):

I would get -70.8y = 0.189

OpenStudy (phi):

add -70.8 to both sides. to do this, write -70.8 on both sides, like this y-70.8 = 0.189x +70.8-70.8 now simplify the right side. what is 70.8 - 70.8 ?

OpenStudy (anonymous):

that is 0

OpenStudy (phi):

yes, so the equation is now y-70.8 = 0.189x or 0.189 x = y-70.8 now divide both sides by 0.189 \[ \frac{0.189 x}{0.189}= \frac{y-70.8}{0.189 } \] the 0.189 divided by itself is 1 so the left side simplifies to 1 times x or just x you now have \[ x= \frac{y-70.8}{0.189 }\] now find x when y is 80.2 can you do that ?

OpenStudy (anonymous):

I get these numbers -374.6031746

OpenStudy (anonymous):

when I divide -70.8 by 0.189

OpenStudy (phi):

you start with the equation \[ x= \frac{y-70.8}{0.189 }\] the problem asks for x when y is 80.2 (it is not clearly written out, but they say y=80.2) so erase the y in the equation, and put in 80.2 in its place can you do that ?

OpenStudy (anonymous):

ok I got that now do I do what we just did just use 80.2 this time

OpenStudy (phi):

erase the y in the equation, and put in 80.2 in its place what is the equation after you do that ?

OpenStudy (anonymous):

80.2 = 0.189x + 70.8

OpenStudy (phi):

ok, but we want to use the equation where we solved for x. \[ x= \frac{y-70.8}{0.189 }\]

OpenStudy (anonymous):

ok so is that the equation

OpenStudy (phi):

yes. the problem said solve the equation for x use this new equation so you want to use this new equation. can you replace y with 80.2 ? what does the equation look like after you do that ?

OpenStudy (anonymous):

0.189x = 80.2 + 70.8

OpenStudy (phi):

replace y with 80.2 in \[ x= \frac{y-70.8}{0.189 }\]

OpenStudy (anonymous):

x= -70.8(80.2)/0.189

OpenStudy (phi):

no, but that is closer. do this \[ x= \frac{\cancel{y}80.2-70.8}{0.189 } \\ x= \frac{80.2-70.8}{0.189 } \]

OpenStudy (anonymous):

I don't have a calculator that dose all that

OpenStudy (anonymous):

that's why I am doing it by hand

OpenStudy (anonymous):

ok I got x =92/0.189

OpenStudy (phi):

do it in steps. what is 80.2 - 70.8 you can type it into google, like this 80.2 - 70.8=

OpenStudy (phi):

80 - 70 is 10 so 80.2 - 70.8 will be close to 10

OpenStudy (anonymous):

ok so is it x = 10/0.189

OpenStudy (phi):

almost 10. type into google this 80.2 - 70.8= what do you get ?

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

9.4 is what I got

OpenStudy (phi):

so now the equation is x= 9.4/0.189 can you find x ?

OpenStudy (phi):

\[ x= \frac{80.2-70.8}{0.189 } \\ x= \frac{9.4}{0.189 } = ?\]

OpenStudy (phi):

you can use google again, if you like: 9.4/0.189=

OpenStudy (anonymous):

ok I got 49. 735449

OpenStudy (phi):

you got x= 49.7 (roughly) the problem says x is the number of years since 1970 what year is that ? I would do 1970+ x to find out.

OpenStudy (anonymous):

so x= 2019

OpenStudy (phi):

to be accurate, x is 49.7 years, and the date is (1970 + x) = 2019.7 the answer to determine in which year the approximate life expectancy will be 80.2 years. is 2019

OpenStudy (anonymous):

ok thank you for the help